the area of a square is 1/4th as large as the area of a triangle. One side of the triangle is 16 inches long, and the altitude to that side is the same length as a side of the square. Find the length of a side of the square.
Answer by mananth(16946)

the area of a square is 1/4th as large as the area of a triangle. One side of the triangle is 16 inches long, and the altitude to that side is the same length as a side of the square. Find the length of a side of the square.
length of square = x
Area of square = x^2
area of triangle = 4x^2

Area of triangle = 1/2 * 16 * x
4x^2 = 8x
4x=8
x=2